OpenStudy (bacon):

from $17.99 to $19.99 what was the percent of increase ?

OpenStudy (anonymous):

11%

OpenStudy (bacon):

may i know how you got that pls?

OpenStudy (bacon):

???????????????????????????????????

OpenStudy (anonymous):

Yes, (1-19.99/17.99)

OpenStudy (anonymous):

?

OpenStudy (bacon):

how did you get the 1?

OpenStudy (anonymous):

11% is right, but that's not the formula

OpenStudy (anonymous):

% increase = (new price - old price) / (old price)

OpenStudy (bacon):

so which one is right?

OpenStudy (anonymous):

(19.99 - 17.99) / 17.99 = 11.12% or rounded to 11%

OpenStudy (bacon):

ok

OpenStudy (anonymous):

@gezimbasha had the right answer, but (1 - 19.99/17.99) = (-18.99/17.99) I think it was just a typo or something.

OpenStudy (bacon):

ok thanks

OpenStudy (anonymous):

Yes, but my formula was this one...\[\frac{19.99}{17.99}-1=\frac{19.99-17.99}{17.99}\]I just swapped the 1 which is still correct but not mathematically legit.. it will just give you -11%.

OpenStudy (bacon):

ok

OpenStudy (bacon):

so would both ways work?

OpenStudy (anonymous):

yeah, its just that I usually just do the first 19.99/17.99 on a calc and subtract from 1, easier for me.

OpenStudy (anonymous):

Oh :) didn't realize what you were doing... that's fine too :)

OpenStudy (anonymous):

as long as you understand how to flip the neg, it's a good calculator short-cut :) @Bacon, if you forget, just try to reason through... % increase is "how much did I increase divided by how much did I start with?"

OpenStudy (bacon):

thanks

OpenStudy (anonymous):

So, if you start with 10 and increase to 15, you increased by 5. 5 over the original 10 means a 50% increase. and so on.

OpenStudy (bacon):

i have another question

OpenStudy (bacon):

i'll open it in something else

OpenStudy (anonymous):

good choice :)